# SRPO: Enhancing Multimodal LLM Reasoning via Reflection-Aware Reinforcement Learning

### v1

@@ 2 Related Works
Reinforcement Learning for LLM Reasoning. Recent advancements in large-scale RL, such as DeepSeek-R1 [4], have demonstrated substantial progress in enhancing complex, human-like Chain-of-Thought (CoT) reasoning by utilizing result-oriented or formatting-specific reward signals. In parallel, several studies, including Open-Reasoner-Zero [29], SimpleRL-Zoo [30], AlphaMed [31], and Logic-RL [32], have explored directly fine-tuning base language models using RL without any supplementary supervised fine-tuning stages. Additionally, methods such as Light-R1 [33] and DeepScaler [34] introduce specially constructed cold-start datasets designed explicitly to encourage detailed step-wise reasoning during initial training phases. Meanwhile, recent analyses [18, 20, 19] have also shed light on intrinsic limitations of purely RL-based reasoning enhancement strategies. Furthermore, complementary approaches such as VAPO [35], DAPO [12], and Dr. GRPO [36] have sought to refine the Group Relative Policy Optimization (GRPO) framework by optimizing reward design and enhancing advantage estimation techniques, thus more effectively promoting deeper reasoning behaviors within language models. In contrast, our work specifically targets multimodal complex reasoning, explicitly emphasizing self-reflection or correction to enhance reasoning performance during both multimodal SFT and RL training phases.
@@ 3.1.1 Self-Reflection SFT data construction
Less is More. To support effective reflection learning, we construct a high-quality dataset generated by advanced MLLMs, containing two types of reflective examples: one for refining correct CoTs by removing redundancy, and another for revising incorrect CoTs through error correction. While prior SFT approaches [24, 25, 26] focus on mimicking correct reasoning, our reflection-oriented SFT explicitly injects reflection knowledge, enabling the model to detect flaws and refine its reasoning. These approaches are complementary—ours enhances reasoning correction and self-improvement. As illustrated in Figure 2, we begin by curating a high-quality subset of $N=10{,}000$ multimodal reasoning samples from three large-scale datasets: LLaVA-CoT [25] (100K), Mulberry [24] (260K), and MathV360K [40]. These samples cover diverse domains including physics, mathematics, and general knowledge. Based on this subset, we construct our self-reflection dataset through two complementary strategies: (1) Refinement of correct CoTs, and (2) Revision of incorrect CoTs. For each sample, we first obtain the initial response generated by the policy model through CoT prompting. Then, using the ground truth answer as guidance, we employ a larger MLLM (e.g., GPT-o4-mini) to generate a self-reflection that either revises flawed reasoning or streamlines correct but verbose outputs. Each final sample thus contains three components: the initial response, the generated self-reflection, and the ground truth answer. In our curated data, approximately 30% of initial responses are correct, while the remaining 70% contain reasoning errors, highlighting the necessity of self-reflection for both wrong solution correction and right question refinement.
@@ 3.1.2 Cold-start Initialization of Policy Model
[…]
@@ 3.2.1 Group Relative Policy Optimization (GRPO)
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})},\quad\text{where}\quad\{r_{i}\}_{i=1}^{G}\quad\text{arerewards from the group.}$. (3)
@@ 3.2.2 Enhanced GRPO with Reflection-Aware Rewards (SRPO)
$R_{\text{task}}=R_{\text{format}}=\begin{cases}0.5,&\text{if format is correct}\\ 0,&\text{otherwise}\end{cases}+R_{\text{accuracy}}=\begin{cases}0.5,&\text{if first solution matches gt label}\\ 0,&\text{otherwise}\end{cases}$ (5)
@@
$I_{\text{eff}}=\begin{cases}0.25,&\text{if reflection keeps a corrected answer},\\ 0.5,&\text{if reflection corrects the wrong answer},\\ 0,&\text{if reflection fails to correct the wrong answer},\\ -0.25&\text{if reflection misconducts the right into wrong answer}.\end{cases}$ (8)
@@ 4.2 Main Results of Benchmarks
Multimodal Mathematical Reasoning. As presented in Table 1, SRPO achieves highly competitive performance on multiple mathematical reasoning benchmarks, even when compared to leading closed-source MLLMs. For instance, on the MathVista benchmark, SRPO obtains a score of 78.5%, trailing the widely acknowledged state-of-the-art model, OpenAI O1, by only 73.9%. Moreover, SRPO consistently outperforms open-source general multimodal baselines by a clear margin. Notably, when compared to state-of-the-art open-source reasoning models such as VL-Rethinker-7B and MM-Eureka-7B, SRPO demonstrates obvious advantages, even on complex, graduate-level reasoning datasets like OlympiadBench. These results strongly validate our claim that explicitly enhancing the model’s self-reflection capabilities during both the SFT and RL stages positively contributes to improved complex reasoning performance.
@@
Table 3: Ablation study of SRPO on RL training data size and self-reflection components.
@@ 4.3 Ablation Study
Effectiveness of Self-Reflection Components. We further investigate individual self-reflection components within SRPO. Table 4.3 shows that removing Self-Reflection SFT notably reduces performance, yet still maintains a 5.1-point advantage over standard GRPO on Physics. Conversely, eliminating Self-Reflection RL yields minimal improvements over Qwen-2.5-VL-7B, indicating that reflection training solely in the SFT stage is insufficient. Hence, explicitly rewarding reflection quality during RL is essential for optimal multimodal reasoning. More analysis about Self-Reflection RL component is illustrated in supplementary part.
@@ 4.4 Reasoning Qualitative Analysis
Self-Reflection in Test Case. Futhermore, figure  3 right part illustrates examples of reasoning outputs during inference. Responses generated by SRPO exhibit clear self-reflection patterns acquired during RL training, actively refining or correcting flawed reasoning steps and answers to improve overall accuracy. Conversely, models trained without self-reflection rarely adjust or rectify incorrect reasoning paths, resulting in persistent reasoning errors and reduced final performance.
@@ A.1 Limitation
Our experiments primarily evaluated the effectiveness of SRPO on dense MLLMs at 7B and 32B scales, without conducting scaling experiments on Mixture-of-Experts (MoE) [62, 63] or diffusion LM [64, 65, 66] architectures. Additionally, the reinforcement learning training data utilized in our experiments were selected exclusively from publicly available multimodal reasoning datasets, without exploration of larger-scale commercial reasoning datasets. Extending our method to MoE-based models and incorporating larger-scale RL training datasets remain promising avenues for future work, potentially leading to even broader improvements in multimodal reasoning capabilities.
@@ Appendix B Appendix
B.1 Training Dataset
@@ B.3 Prompt Template
B.4 Generated Samples Analysis
@@
Generated Samples in Real Test cases of GRPO (Sample 3) Question: A farmer wants to build a fence around the entire perimeter of his land, as shown in the diagram. The fencing costs £37 per metre. The lengths of $EF$ and $CD$ are $x$ metres and $y$ metres respectively. At £37 per metre, how much will it cost him to build the fence along the entire perimeter?  Ground truth: 777[Uncaptioned image] <think>
@@
The initial reasoning misidentified the structure of the perimeter by incorrectly including internal segments such as $BG$, $GE$, and $EB$. These are auxiliary construction lines used solely to form right triangles for computing the diagonal lengths $EF$ and $CD$, not actual parts of the outer boundary of the land that require fencing. This led to an overestimation of the total perimeter.
To correct the reasoning, one must focus strictly on the visual geometry and trace only the external contour of the land—specifically, the true fencing path comprises segments $AB$, $AF$, $EF$, $ED$, $DC$, and $CA$. The diagonal lengths $EF$ and $CD$ should be computed using the Pythagorean theorem based on right triangle constructions derived from the diagram. This corrected understanding yields the true perimeter and thus the accurate total fencing cost.
@@
$AB=5,\ AF=1,\ FE=x,\ ED=2,\ DC=y,\ CB=2.$ Therefore the total perimeter is $5+1+x+2+y+2=10+(x+y).$ From the diagram’s length proportions (and the single “11 m” label spanning the two slanted edges), we see directly that $x+y=11.$ Hence the full perimeter is $10+11=21$ m, and the fencing cost is $37\times 21=\boxed{777}.$ </think>
@@
[…]
@@
[…]

### v3

@@ 2 Related Works
Reinforcement Learning for LLM Reasoning. Recent advancements in large-scale RL, such as DeepSeek-R1 [4], have demonstrated substantial progress in enhancing complex, human-like Chain-of-Thought (CoT) reasoning by utilizing result-oriented or formatting-specific reward signals. In parallel, several studies, including Open-Reasoner-Zero [29], SimpleRL-Zoo [30], AlphaMed [31], and Logic-RL [32], have explored directly fine-tuning base language models using RL without any supplementary supervised fine-tuning stages. Additionally, methods such as Light-R1 [33] and DeepScaler [34] introduce specially constructed cold-start datasets designed explicitly to encourage detailed step-wise reasoning during initial training phases. Meanwhile, recent analyses [18, 20, 19] have also shed light on intrinsic limitations of purely RL-based reasoning enhancement strategies. Furthermore, complementary approaches such as VAPO [35], DAPO [12], PTA-GRPO [36], and Dr. GRPO [37] have sought to refine the Group Relative Policy Optimization (GRPO) framework by optimizing reward design and enhancing advantage estimation techniques, thus more effectively promoting deeper reasoning behaviors within language models. In contrast, our work specifically targets multimodal complex reasoning, explicitly emphasizing self-reflection or correction to enhance reasoning performance during both multimodal SFT and RL training phases.
@@ 3.1.1 Self-Reflection SFT data construction
Less is More. To support effective reflection learning, we construct a high-quality dataset generated by advanced MLLMs, containing two types of reflective examples: one for refining correct CoTs by removing redundancy, and another for revising incorrect CoTs through error correction. While prior SFT approaches [24, 25, 26] focus on mimicking correct reasoning, our reflection-oriented SFT explicitly injects reflection knowledge, enabling the model to detect flaws and refine its reasoning. These approaches are complementary—ours enhances reasoning correction and self-improvement. As illustrated in Figure 2, we begin by curating a high-quality subset of N=10, [^0] N=10{,}000 multimodal reasoning samples from three large-scale datasets: LLaVA-CoT [25] (100K), Mulberry [24] (260K), and MathV360K [41]. These samples cover diverse domains including physics, mathematics, and general knowledge. Based on this subset, we construct our self-reflection dataset through two complementary strategies: (1) Refinement of correct CoTs, and (2) Revision of incorrect CoTs. For each sample, we first obtain the initial response generated by the policy model through CoT prompting. Then, using the ground truth answer as guidance, we employ a larger MLLM (e.g., GPT-o4-mini) to generate a self-reflection that either revises flawed reasoning or streamlines correct but verbose outputs. Each final sample thus contains three components: the initial response, the generated self-reflection, and the ground truth answer. In our curated data, approximately 30% of initial responses are correct, while the remaining 70% contain reasoning errors, highlighting the necessity of self-reflection for both wrong solution correction and right question refinement.
@@ 3.1.2 Cold-start Initialization of Policy Model
Here, $a_{1}$ is the policy model’s initial response, $\langle\texttt{reflection}\rangle$$\langle\texttt{/reflection}\rangle$ denotes the reflection generated by a large LLM, and $a_{2}$ is the ground truth answer. Given the input prompt $q$ and the policy model $\pi_{\text{initial}}$, the objective is twofold: (1) to train the model to revise $a_{1}$ toward $a_{2}$ using the reflection $\langle\texttt{reflection}\rangle$, and (2) to leverage the reasoning and knowledge embedded in $a_{2}$ to guide future predictions. This reflection-driven learning process equips the policy model with self-correction capabilities and improves alignment with correct reasoning trajectories.
@@ 3.2.1 Group Relative Policy Optimization (GRPO)
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})},\quad\text{where}\quad\{r_{i}\}_{i=1}^{G}\quad\text{are rewards from the group.}$ (3)
@@ 3.2.2 Enhanced GRPO with Reflection-Aware Rewards (SRPO)
[…]
@@
[…]
@@ 4.2 Main Results of Benchmarks
Multimodal Mathematical Reasoning. As presented in Table 1, SRPO achieves highly competitive performance on multiple mathematical reasoning benchmarks, even when compared to leading closed-source MLLMs. For instance, on the MathVista benchmark, SRPO obtains a score of 78.5%, trailing the widely acknowledged state-of-the-art model, OpenAI GPT-o1, by only 73.9%. Moreover, SRPO consistently outperforms open-source general multimodal baselines by a clear margin. Notably, when compared to state-of-the-art open-source reasoning models such as VL-Rethinker-7B and MM-Eureka-7B, SRPO demonstrates obvious advantages, even on complex, graduate-level reasoning datasets like OlympiadBench. These results strongly validate our claim that explicitly enhancing the model’s self-reflection capabilities during both the SFT and RL stages positively contributes to improved complex reasoning performance.
@@
Table 3: Ablation study of SRPO-7B on RL training data size and self-reflection components.
@@
- no Length Reward ($f_{\text{len}}(\cdot)$) 37K 75.3 56.2 32.4 41.7 60.1 53.1
- no Effectiveness Reward ($I_{\text{eff}}$) 37K 73.9 54.7 31.6 40.9 58.8 52.0
@@ 4.3 Ablation Study
Effectiveness of Self-Reflection. We further investigate individual self-reflection components within SRPO. Table 4.3 shows that removing Self-Reflection SFT notably reduces performance, yet still maintains a 5.1-point advantage over standard GRPO on Physics. Conversely, eliminating Self-Reflection RL yields minimal improvements over Qwen-2.5-VL-7B, indicating that reflection training solely in the SFT stage is insufficient. Hence, explicitly incentivizing reflection behavior during RL is essential for ehancing multimodal reasoning.
Effectiveness of Reflection-Aware RL Components. We also observe that omitting any specific Self-Reflection RL component can degrade performance, especially when the Effectiveness Reward ($I_{\text{eff}}$) is removed, resulting in a drop in average performance from 53.5 to 52.0. It indicates that the model critically relies on reward signals that explicitly evaluate the quality of reflective responses to achieve optimal reasoning. Similarly, reducing the Length Reward ($f_{\text{len}}$) also leads to a decline in reasoning performance, suggesting that overly redundant thinking steps can interfere with the model’s accurate reasoning.
Table 4: Comparison between SRPO and GRPO with 2-Step Thinking.
Methods MathVista MathVerse MathVision MMMU-Pro Physics Avg.
$\star$ (Ours - SRPO-7B) 75.8 55.8 32.9 42.3 60.6 53.5
GRPO 72.3 52.9 30.3 39.9 53.5 49.8
GRPO + 2-Step Thinking 73.5 53.6 30.6 40.3 53.6 50.3
Comparative Study of Reflective and Two-Step Thinking. To assess the effectiveness of our proposed reflection pattern—comprising thinking, reflection, and rethinking—we compare it against a GRPO-based 2-step thinking paradigm, where the model generates two consecutive <think>...</think> steps. Training uses the task reward from Equation 5 and a relational reward inspired by Equation 8 that captures consistency between the two steps. As shown in Table 4, GRPO + 2-Step Thinking offers no significant gains over vanilla GRPO, except on MathVerse. In contrast, SRPO’s explicit reflection on prior thinking substantially improves reasoning, underscoring the importance of combining Self-Reflection SFT with Reflection-aware RL.
Disentangling the Effect of Reflection Format vs. Teacher Distillation To examine whether the SFT improvement stems from teacher distillation or the proposed reflection structure, we conduct controlled experiments on the 7B model. All models share the same teacher (GPT-4-mini), dataset size (10k), optimizer, and number of training epochs. The only differences lie in the inclusion of <reflection> segments and whether reflection-aware RL is applied. The results are shown in Table 5. Adding explicit reflection segments in SFT yields an average gain of +1.3 over plain CoT SFT, indicating that structured reflection rather than teacher distillation contributes to better reasoning performance. When both models undergo RL, SRPO achieves an additional +2.1 gain by aligning rewards with reflection effectiveness, which highlights the benefit of reflection-aware reward design.
Table 5: Ablation of Self-Reflection SFT on the 7B model.
Methods MathVista MathVerse MathVision MMMU-Pro Physics Avg.
Qwen-2.5-VL-7B 68.2 46.3 25.1 36.9 45.4 44.4
Plain-CoT SFT 69.1 47.2 26.4 37.2 47.6 45.3
Reflection-SFT (ours) 70.3 48.2 27.2 38.7 48.5 46.6
Plain-CoT SFT + GRPO 73.6 54.2 30.6 40.6 58.0 51.4
Reflection-SFT + Reflection-RL 75.8 55.8 32.9 42.3 60.6 53.5
@@ 4.4 Reasoning Qualitative Analysis
Self-Reflection in Test Case. Futhermore, figure 3 right part illustrates examples of reasoning outputs during inference. Responses generated by SRPO exhibit clear self-reflection patterns acquired during RL training, actively refining or correcting flawed reasoning steps and answers to improve overall accuracy. Conversely, models trained without self-reflection rarely adjust or rectify incorrect reasoning paths, resulting in persistent reasoning errors and reduced final performance.
@@ 6 Acknowledgment
7 Contributions
Project Lead  [^1] Work done during internship at ByteDance Seed.
Zhongwei Wan2
Core Contributors
Zhongwei Wan2, Zhihao Dou3
Contributors
Che Liu4, Yu Zhang11, Dongfei Cui5, Qinjian Zhao6, Hui Shen7, Jing Xiong10, Yi Xin12, Yifan Jiang8, Chaofan Tao10, Yangfan He9
Supervisors
Mi Zhang2, Shen Yan1
Affiliation
1ByteDance Seed
2The Ohio State University
3Case Western Reserve University
4Imperial College London
5Duke University
6Kean University
7University of Michigan
8University of Southern California
9University of Minnesota
10The University of Hong Kong
11Tongji University
12Nanjing University
@@ A.1 Limitation
Our experiments primarily evaluated the effectiveness of SRPO on dense MLLMs at 7B and 32B scales, without conducting scaling experiments on Mixture-of-Experts (MoE) [63, 64] or diffusion LM [65, 66, 67] architectures. Additionally, the reinforcement learning training data utilized in our experiments were selected exclusively from publicly available multimodal reasoning datasets, without exploration of larger-scale commercial reasoning datasets. Extending our method to MoE-based models, incorporating larger-scale RL training datasets, and combining with efficient inference [68, 69, 70, 71, 72, 73, 74, 75, 76, 77] or other decoding settings [78, 79] remain promising avenues for future work, potentially leading to even broader improvements in multimodal reasoning capabilities.
@@ Appendix B Appendix
B.1 Further Experiment
Sensitivity to the Reflection Length Reward Coefficient $\alpha$. To evaluate the robustness of SRPO to the reflection-length reward coefficient, we additionally trained models with $\alpha=0.05$ and $\alpha=0.3$, using the same setup as in the main experiments. The results are shown in Table 6. All experiments were performed under identical optimization and data settings, with only the length-reward coefficient varied. Results indicate that $\alpha=0.05$ performs similarly to $\alpha=0.1$, while $\alpha=0.3$ slightly degrades accuracy. This suggests that over-emphasizing length reward weakens the reflection signal, whereas values near $0.1$ offer a balanced trade-off, helping to control response length and stabilize training with limited impact on accuracy.
Table 6: Sensitivity of SRPO to the reflection length reward coefficient $\alpha$.
Ratio MathVerse MMMU-Pro Physics Avg.
$\alpha$=0.05 55.7 41.9 61.1 52.9
$\alpha$=0.1 (Ours) 55.8 42.3 60.6 53.5
$\alpha$=0.3 55.1 42.0 59.7 52.2
Qualitative & quantitative reflection analysis. To quantitatively assess reflection effectiveness, we conducted both human and LLM-based evaluations on SRPO-7B’s generated reflections.
a. Human Expert Evaluation. We randomly sampled 100 MathVista test questions (answered by SRPO-7B). Two senior PhD students specializing in NLP and LLMs independently rated each reflection on a 0–3 scale: 3 = highly effective, 2 = partially effective, 1 = redundant, and 0 = detrimental. We also measured the Wrong Answer Fix Rate, i.e., how often an initially incorrect answer was corrected after reflection. The results are summarized in Table 7.
Table 7: Human Expert Evaluation of Reflection Quality.
Metric Human Expert 1 Human Expert 2
Effective Reflection Rate (score $\geq$ 2) 73% 69%
Redundancy Rate (score = 1) 9% 11%
Detrimental Rate (score = 0) 3% 1%
Wrong Answer Fix Rate 39% 39%
Seventy percent of reflections were judged to be effective. Among 100 questions, the initial solution was incorrect in 33 cases, of which 13 were corrected after reflection, resulting in a 39% wrong-answer fix rate.
b. LLM-as-a-Judge (GPT-4o) Evaluation. Each reflection was further scored by GPT-4o on four 0–5 dimensions—logical flaws, missing assumptions, clarity, and actionable suggestions—following the prompt format described in Appendix B.3. The average reflection quality is shown in Table 8.
Table 8: LLM-as-a-Judge Evaluation (GPT-4o).
Model Logic Missing Clarity Suggestions Avg. Quality
SRPO-7B (Ours) 4.1 3.9 4.0 3.6 3.9
These results confirm that SRPO produces high-quality reflections from both human and LLM perspectives, validating the effectiveness of its self-reflection design.
@@ B.1 Further Experiment
Training and Inference Efficiency
We report the measured wall-clock time for each SRPO training stage under both 7B and 32B settings using the OpenRLHF framework. All settings share the same data size and optimization schedule as in the main experiments; only the model size and stage differ. Table 9 summarizes training efficiency, and Table 10 summarizes inference efficiency on the MathVista test-mini split ( 1k samples) using a single H100-80G GPU with vLLM. SRPO requires only modest additional wall-time compared with GRPO, while delivering substantially better reasoning quality in the main results.
Table 9: Training efficiency overview. Wall-clock time for each stage under 7B and 32B settings.
Training Stage GPUs Wall-Time Notes
Reflection-SFT (7B) $8\timesH$100 3.5 hours 10k samples, 1 epoch
Reflection-RL (7B, SRPO) $8\timesH$100 31.2 hours 37k samples, 500 steps
GRPO-RL (7B) $8\timesH$100 25.8 hours 37k samples, 500 steps
Reflection-SFT (32B) $4\times 8\timesH$100 4.7 hours 10k samples, 1 epoch
Reflection-RL (32B, SRPO) $4\times 8\timesH$100 45.1 hours 37k samples, 500 steps
Table 10: Inference efficiency overview. Latency on MathVista test-mini using a single H100-80G with vLLM.
Inference Stage GPUs Wall-Time Notes
GRPO-RL (7B) $1\timesH$100 45.6 min 1k test samples, avg response = 355.4 tokens
SRPO (7B) $1\timesH$100 60.5 min 1k test samples, avg response = 502.3 tokens
B.2 Training Dataset
@@ B.4 Prompt Template
B.5 Generated Samples Analysis
@@
Generated Samples in Real Test cases of GRPO (Sample 3) Question: A farmer wants to build a fence around the entire perimeter of his land, as shown in the diagram. The fencing costs £37 per metre. The lengths of EFEF and $CD$ are $x$ metres and $y$ metres respectively. At £37 per metre, how much will it cost him to build the fence along the entire perimeter?    Ground truth: 777[Uncaptioned image] <think>
@@
The initial reasoning misidentified the structure of the perimeter by incorrectly including internal segments such as BGBG, GEGE, and EBEB. These are auxiliary construction lines used solely to form right triangles for computing the diagonal lengths EFEF and $CD$, not actual parts of the outer boundary of the land that require fencing. This led to an overestimation of the total perimeter.
To correct the reasoning, one must focus strictly on the visual geometry and trace only the external contour of the land—specifically, the true fencing path comprises segments ABAB, AFAF, EFEF, EDED, DCDC, and CACA. The diagonal lengths EFEF and CDCD should be computed using the Pythagorean theorem based on right triangle constructions derived from the diagram. This corrected understanding yields the true perimeter and thus the accurate total fencing cost.
@@
AB=5,AF=1,FE=x,ED=2,DC=y,CB=2.AB=5,\ AF=1,\ FE=x,\ ED=2,\ DC=y,\ CB=2. Therefore the total perimeter is $5+1+x+2+y+2=10+(x+y).$ From the diagram’s length proportions (and the single “11 m” label spanning the two slanted edges), we see directly that $x+y=11.$ Hence the full perimeter is $10+11=21$ m, and the fencing cost is $37\times 21=\boxed{777}.$ </think>
@@
B, C, and D are chemically consistent. Therefore, the incorrect statement is $\boxed{\text{A}}$.
@@
From periodic positions and common valences: M = Na, W = Mg, Y = Al, Z = P. Only statement A is chemically incorrect. Answer: $\boxed{\text{A}}$.